If the current (I) in a long solenoid is doubled and number of turns per unit length (n) are halved, find the effect on the magnetic field (i) inside and (ii) outside the solenoid.


1
Expert's answer
2022-03-21T12:47:22-0400

Solenoid magnetic field

"B=\frac{\mu_0Ni}{L}"

Inside


"B=\mu_0 ni"

"B'=\mu_0n'i'"

"B'=\mu_0\times\frac{n}{2}\times2i"

"\frac{B'}{B}=\frac{1}{1}"

"B'=B"

Part(b)

Out side

"B_o=0T"

Outside magnetic field of solenoid at any point zero

"B_o=0T"